OpenStudy (anonymous):

The average number of vehicles waiting in line to enter a sports arena parking lot is modeled by the function w(x)=x^2/2(1-x), where x is a number between 0 and 1 known as the traffic intensity. Find the average number of vehicles waiting if the traffic intensity is 0.92.

OpenStudy (ash2326):

Average no. of vehicles w(x)=x^2/2(1-x) x=0.92 w(x)=0.92^2/2(1-0.92) =0.92^2/0.16 =5.29 Average 5.29 vehicles are waiting

OpenStudy (anonymous):

thanks man